[figure: 9.7 cm]

MAGNETIC PROPERTIES OF DyMn_{2} SINGLE CRYSTAL

E. Talik^{a}, M. Kulpa^{a}, T. Mydlarz^{b}, J. Kusz^{a}, H. Böhm^{c}, J. Kubacki^{a} aInstitute of Physics, University of Silesia, Uniwersytecka 4, 40-007 Katowice, Poland

bInstitute of Low Temperatures and High Fields, Gajowicka 95, 53-529 Wrocław, Poland cInstitut für Geowissenschaften der Universität, 55099 Mainz, Germany

For RMn_{2} compounds the instability of Mn magnetic moment is observed. DyMn_{2} single crystals were obtained by the Czochralski method from the levitated melt. The sample’s quality was checked by the X-ray powder diffraction using Siemens D-5000 diffractometer with Cu K_{α} radiation and Berg-Barrett topography with Fe K_{α} radiation. There were also taken the XPS spectra on the single crystalline sample. The electrical resistivity and ac magnetic susceptibility along growth [111] direction was measured within temperature range 4.2-300 K showing the change in slope at T_{N}=36 K. Lattice parameters temperature dependence was measured from T =10 K to the room temperature. Magnetic measurements in the International Laboratory of High Fields and Low Temperatures in Wrocław in the temperature range 4.2-300 K are in progress. The measurements of susceptibility χ_{ac} and χ_{dc} in the wide temperature range are in progress.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : E. Talik

Address for correspondence : University of Silesia Institute of Physics Uniwersytecka 4 40-007 Katowice

Email address : talik@us.edu.pl